OpenStudy (anonymous):

Choose the correct simplification of the expression (8x2)3. 24x6 512x6 512x5 24x5

OpenStudy (unklerhaukus):

\[(8x^2)^3=8^3x^{2\times3} \]

OpenStudy (anonymous):

i dont understand?

OpenStudy (unklerhaukus):

\[(a^xb^y)^n=a^{xn}b^{yn}\]

OpenStudy (anonymous):

24*6

OpenStudy (anonymous):

8 cubed isnt 24 it beh the second one

OpenStudy (anonymous):

@UnkleRhaukus i know that the outside exponent multiply's into the equation

OpenStudy (unklerhaukus):

what dont you understand then?

OpenStudy (anonymous):

i think its b...

OpenStudy (anonymous):

like how to multiply them after they are put into the equation

OpenStudy (anonymous):

abd with this one would it look like this "8x^23

OpenStudy (anonymous):

and not abd